OpenStudy (anonymous):

An electron is placed in an electric field and is allowed to accelerate for a distance of 1 m. If the strength of the electric field is E = 200 V/m then what is the final velocity of the electron, assuming it starts from rest?

OpenStudy (ujjwal):

Force on electron, F=eE.......(1) where e is electronic charge and E is electric field also, gain in KE = work done or, 1/2 mv^2 = FS or, 1/2 mv^2 = eES \[or, v=\sqrt{(2eES)/m}\] where m is mass of electron S is distance traveled by electron v is final velocity attained your answer will be 8.38*10^6 m/s (approx)

OpenStudy (anonymous):

thanks, the answer is 8.4 × 106 m/s :)

OpenStudy (anonymous):

10^6*